# Leveraging Atom Loss Errors in Fault Tolerant Quantum Algorithms

### v2

@@ B.2 Approximate MLE Decoding
In this section, we describe our approach for constructing a decoder to handle losses by approximating the MLE decoding introduced in the previous section.
We utilize the concept of qubit lifecycles: the number of potential locations where a physical qubit may be lost—due to single-qubit gates, multi-qubit gates, initialization, idling, measurement, or other processes. Given a heralded loss upon measurement, the decoder identifies a loss occurring in one of these locations, each characterized by its specific probability of loss.
Consider a scenario where, at the end of qubit $i$’s lifecycle $C_{i}$, a loss is detected, with $L_{ij}$ representing potential loss events, each associated with a probability $p_{ij}$. For each potential loss event $L_{ij}$, we simulate the effect of the loss by constructing a loss circuit. This involves removing subsequent gates acting on the lost qubit and modeling the measurement of the lost qubit as random. Using Stim’s gauge detector infrastructure [53], we generate the decoding graph for each specific loss event, denoted as $\text{DEM}_{ij}$.
The decoding graph is a collection of hyperedges and their associated probabilities. Each error mechanism corresponds to a hyperedge that connects the detectors it triggers. Detectors are defined as the products of stabilizer measurements and are expected to be $+1$ in the absence of errors.
Each decoding graph $\text{DEM}_{ij}$ is a collection of error probabilities $p_{n}$ and corresponding detector configurations $D_{n}$:
@@
where $D_{n}=[D_{n}(0),D_{n}(1),\cdots]$ represents the list of detectors triggered by the error event. Repeating this process for all potential loss events $L_{ij}$ within a lifecycle $C_{i}$ results in a set of decoding graphs $\text{DEM}_{ij}$ and their associated probabilities $p_{ij}$.
To compute the final decoding graph for a lifecycle $C_{i}$, we sum over all decoding graphs for the potential loss events:
$\text{DEM}_{i}=\sum_{j}p_{ij}\cdot\text{DEM}_{ij}.$ (4)
This summation is straightforward because only one of these events can occur.
@@
To address this, we explore a decoder that evaluates each lifecycle independently and averages over them. Additionally, we introduce the option to include a single combination of losses across different lifecycles. Specifically, we consider the first potential loss locations $L_{i1}$ for all lossy lifecycles $L_{i}$ and generate a corresponding decoding graph, $\text{DEM}_{\text{first comb}}$.
The decoding graphs for individual potential loss locations $L_{ij}$ are generated in a pre-processing step, independent of any specific error model or probabilities of Pauli and loss errors. During real-time decoding, these graphs are summed using the specific error model probabilities $p_{ij}$. Note that we separately account for loss errors and Pauli errors. Lossless circuits (without losses) are used to efficiently generate the decoding graph for Pauli errors, denoted $\text{DEM}_{\text{Pauli}}$.
For a specific shot with a heralded loss pattern using SSR and a given error model, the final decoding graph is computed by summing: 1. Lossy lifecycle decoding graphs $\text{DEM}_{i}$, 2. The Pauli decoding graph $\text{DEM}_{\text{Pauli}}$, and 3. The first potential loss combination decoding graph $\text{DEM}_{\text{first comb}}$.
The final decoding graph is given by:
@@
where $\omega$ is a combination weight determined based on the analysis in Appendix Section B.2.B.2.1.
@@
Finally, given the final decoding graph, any known decoder can be used to decode the data for a specific shot, such as MWPM, MLE, or other methods.
@@ B.2.1 Adjusting Loss Combination Weight
In this section, we explore the influence of the weight of the combination, $\omega$, on the efficacy of the delayed erasure decoder, which considers both independent error events and the first combination of potential loss events. As described earlier, the decoder’s decision-making process involves summing different decoding graphs (see Eq. 5). To incorporate the impact of individual loss events, the first loss combination event is scaled by $\omega$ along with the standard Pauli error decoding graph.
To benchmark the performance of our decoder and determine the optimal value of $\omega$, we performed circuit-level simulations on a memory logical qubit with multiple rounds of conventional syndrome extraction (SE). We varied $\omega$ between 0 and 1. For $\omega=0$, each lifecycle’s decoding graph is considered independently, without including an additional decoding graph for the first combination.

### v3

@@ B.2 Approximate MLE Decoding
Figure 9: Overview of the approximate MLE decoding procedure. (a) For a given qubit lifecycle, we identify potential loss locations (red crosses) and simulate modified circuits for each, generating a set of decoding hypergraphs (DEMij). These include updated stabilizer syndromes (detectors) and correlated error edges resulting from the truncated loss circuits. (b) When multiple qubit losses are present in a given shot, we compute the decoding hypergraph for each lifecycle independently (left), optionally include a combined loss decoding hypergraph (right), and sum with the standard Pauli decoding hypergraph to generate the final decoding hypergraph (DEMfinal), according to Eq. 6. This forms the input to the decoder.
In this section, we describe our practical decoding strategy that approximates the MLE decoding introduced above. The goal is to capture the key features of loss-induced error correlations, particularly the way loss truncates gate sequences and alters detector structure, without solving the full combinatorial problem.
As stated in the main text, our goal is to automatically construct a decoding hypergraph based on the observed loss, which serves as the input to any general decoder for the QEC code analyzed. Our general approach relies on precomputing the decoding hypergraph for each possible loss location within a qubit’s lifecycle. Given a heralded loss upon measurement, we infer that the loss occurred at some earlier location in the lifecycle, and assign a probability to each possible location. Given a specific experiment with a given loss pattern, we can sum over the relevant precomputed lifecycle decoding hypergraphs, as described below. Figure 9 illustrates this procedure, showing how the loss locations in each lifecycle contribute to the overall decoding hypergraph for a given shot.
We now formalize our approach. Let $C_{i}$ denote the lifecycle of qubit $i$, and let $\{L_{ij}\}$ denote the set of possible loss locations for that qubit, each associated with a probability $p_{ij}$. For each possible loss event $L_{ij}$, we simulate the loss circuit, a modified circuit generated by removing all gates after the loss and setting the qubit’s final measurement to a uniformly random outcome. Using Stim’s gauge detector infrastructure [53], we extract the decoding hypergraph denoted $\text{DEM}_{ij}$, which consists of a set of hyperedges (error events) and their associated detector configurations.
Each decoding hypergraph $\text{DEM}_{ij}$ is a matrix, a collection of error probabilities $p_{n}$ and corresponding detector configurations $D_{n}$:
@@
where the columns $D_{n}=[D_{n}(0),D_{n}(1),\cdots]$ represent the list of detectors triggered by the error event in each row. Repeating this process for all potential loss events $L_{ij}$ within a lifecycle $C_{i}$ results in a set of decoding hypergraphs $\text{DEM}_{ij}$ and their associated probabilities $p_{ij}$. This is done in a pre-processing stage.
To compute the final decoding hypergraphs for a lifecycle $C_{i}$ of a lost qubit $i$, we sum over all decoding hypergraphs in the lifecycle:
$\text{DEM}_{i}=\sum_{j}\tilde{p}_{ij}\cdot\text{DEM}_{ij}.$ (5)
with normalized weights $\tilde{p}_{ij}=\Pr(L_{ij}\mid\text{flag in }C_{i})$; this is an approximation that assumes at most one loss per lifecycle and neglects cross-lifecycle combinations.
@@
To address this, we explore a decoder that evaluates each lifecycle $C_{i}$ independently and averages over them ($\text{DEM}_{i}$). Additionally, we introduce the option to include a single combination of losses across different lifecycles. Specifically, we consider the first potential loss locations $L_{i1}$ for all lossy lifecycles $L_{i}$ and generate a corresponding decoding hypergraph, $\text{DEM}_{\text{first comb}}$.
The decoding hypergraphs for individual potential loss locations $L_{ij}$ are generated in a pre-processing step, independent of any specific error model or probabilities of Pauli and loss errors. During real-time decoding, these hypergraphs are summed using the specific error model probabilities $\tilde{p}_{ij}$. Note that we separately account for loss errors and Pauli errors. Lossless circuits (without losses) are used to efficiently generate the decoding hypergraph for Pauli errors, denoted $\text{DEM}_{\text{Pauli}}$.
For a specific shot with a heralded loss pattern using SSR and a given error model, the final decoding hypergraph is computed by summing: 1. Lossy lifecycle decoding hypergraphs $\text{DEM}_{i}$, 2. The Pauli decoding hypergraph $\text{DEM}_{\text{Pauli}}$, and 3. The first potential loss combination decoding hypergraph $\text{DEM}_{\text{first comb}}$.
The final decoding hypergraph is given by:
@@
where $\omega$ is a combination weight determined based on the analysis in Appendix Section B.B.2.B.2.1.
@@
Once $\text{DEM}_{\text{final}}$ is constructed for a given shot, any decoder that accepts a hypergraph-based detector error model may be used—such as MLE, MWPM, belief propagation or others, depending on the QEC code in use.
@@ B.2.1 Adjusting Loss Combination Weight
In this section, we explore the influence of the weight of the combination, $\omega$, on the efficacy of the delayed erasure decoder, which considers both independent error events and the first combination of potential loss events. As described earlier, the decoder’s decision-making process involves summing different decoding hypergraphs (see Eq. 6). To incorporate the impact of individual loss events, the first loss combination event is scaled by $\omega$ along with the standard Pauli error decoding hypergraph.
To benchmark the performance of our decoder and determine the optimal value of $\omega$, we performed circuit-level simulations on a memory logical qubit with multiple rounds of conventional syndrome extraction (SE). We varied $\omega$ between 0 and 1. For $\omega=0$, each lifecycle’s decoding hypergraph is considered independently, without including an additional decoding hypergraph for the first combination.